. A base line was measured to be 150 m long with a tape at a field temperature of 27°C, the applied pull being 14 kg. The tape was standardised at a temperature of 15°C with a pull of 8 kg. If the designated length of the tape is 20 m, weight of 1 cm of tape = 7.86 g, total weight of tape = 0.8 kg, E = 2.109 x 106 kg/cm² and coefficient of expansion of tape per degree Celsius 11.2 x 10°, find the true length of the line. (Ans. 150.0081 m)


1
Expert's answer
2021-09-28T00:54:43-0400

Temperature correction per tape length=

"( Ct\n ) = \u03b1(Tm\u2212T0)L"

"=3.5\u00d710\u22126(34\u201325)\u00d730"

"=9.45 \u00d7 10^{-4} (+VE)"


To find pull correction per tape length(Pull correction per tape length)

"= \frac{p_{m-P_0 } L}{AE}=\frac{(18-10)30}{(0.025 × 2.1× 10^6 ) } = 150.008 \space m"